OpenStudy (anonymous):

solve the equation for a 2sqrt a-1= 4 sqrt a

OpenStudy (anonymous):

\[2\sqrt{a-1}=4\sqrt{a}\] this is how the problem reads

OpenStudy (anonymous):

\[\sqrt[2]{a-1}=\sqrt[4]{a}\] is the question

OpenStudy (anonymous):

yes and i need to figure out what a is and i have no idea how

OpenStudy (anonymous):